OpenStudy (arbyj):

helper will get a medal! Write the standard form of the line that passes through the given points. Include your work in your final answer. (1, 5) and (-2, 3) ive already found the slope

OpenStudy (welshfella):

use the point-slope form of the equation of a line y-y1 = m(x-x1) (m =slope and (x1,y) is a point on the line)

OpenStudy (welshfella):

use any of the given points

OpenStudy (arbyj):

just to make sure the slope is 2/3, righ?

OpenStudy (welshfella):

finally rearrange to standard form ax + by = c

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

so if we use the point (1,5) we get y - 5 = (2/3)( x- 1) the next step is the divide through by 3 to get rd of the fraction.

OpenStudy (welshfella):

can yu continue ?

OpenStudy (arbyj):

would it -2x+3y=4? @welshfella

OpenStudy (welshfella):

y - 5 = 2/3 (x - 1) * by 3 3y - 15 = 2(x - 1) 3y - 15 = 2x - 2 -2x + 3y = -2 + 15 -2x + 3y = 13

OpenStudy (welshfella):

lloks like you did not multiply the -5 by 3

OpenStudy (arbyj):

oh okay, thank you!